OpenStudy (anonymous):

Evaluate: 9-3x2+4= step by step A.6 B.7 C.15 D.16

OpenStudy (anonymous):

is that = 0?

OpenStudy (anonymous):

if so then -3x^2 = -4-9 ===> -3x^2=-13 ====> x^2 = -13/-3 ===> x = sqrt(13/3)

OpenStudy (anonymous):

I have to find the answer I think it is B but I don't know if I am correct.

OpenStudy (anonymous):

do you have a specific value for x?

OpenStudy (anonymous):

Its a multiplication sign do you use x or * for that?

OpenStudy (anonymous):

I keep trying to solve the problem but I always get -1 but I think the answer is b

OpenStudy (anonymous):

oh ok hahaha next time just use * its alot clearer

OpenStudy (anonymous):

9-3*2 +4 = 9 - 6 + 4 = you always do left to right = 3 + 4 = 7

OpenStudy (phi):

use order of operations... multiplication before add so 9-3x2+4 means do 3*2 = 6 so you have 9 - 6 +4 now add in order 9+ -6 (or 9-6 which means the same thing) then add 4

OpenStudy (anonymous):

Oh okay now I understand but why subtract 9-6 first the add 3+4? isn't addition first then subtraction?

OpenStudy (phi):

when you have 9 - 6 +4 you would normally do it left to right if you write the subtraction as adding a negative number, like this 9 + -6 + 4 then you are adding 3 numbers, and as you know you can switch the order when adding 1+2 is the same as 2+1, or 9 + -6 +4 is the same as 4+9 + -6 I always think of subtraction as adding a negative number...

OpenStudy (anonymous):

Ooh THANK YOU I always would get this type of problem wrong :D

OpenStudy (anonymous):

subtracting and adding are the same thing, and multiplying and dividing are the same thing so you if you have that then just go left to right

OpenStudy (phi):

another way to think of your problem 9-3x2+4 is 9 + -3*2 + 4 (I switched to * because in algebra x stands for "unknown") now it is still order of operations: multiply before add, but we do -3*2 = -6 and get 9 + -6 + 4 which gives 7 , just like before

OpenStudy (phi):

I hope that doesn't confuse you